Hamming Code: m = 1010 (4-bit message); odd-parity scheme; Compute Hamming Code on the sender’s side. On the receiver’s side, detect and verify a 7th bit error-position in the transmitted data.
Answer #1

Length of sequence-4

Number of parity bit required-4

length of code word=8

parity sceheme to be used,

P0=D0+D1+D3

P1=D0+D2+D3

P2=D1+D2+D3

P3=D0+D1+D2+P0+P1+P2

D3 D2 D1 D0 P3 P2 P1 P0
1 1 1 1 1 1 1 1
1 1 1 0 0 1 0 0
1 1 0 1 0 0 1 0
1 0 1 1 0 0 0 1

So,the encoded value would be 10101010

Error in 7th bit detection means checking P1 collumn for correct value..As we can see in table its not showing an error.

D3 D2 D1 D0 P3 P2 P1 P0
1 1 1 1 1 1 1 1
1 1 1 0 0 1 0 0
1 1 0 1 0 0 1 0
1 0 1 1 0 0 0 1

If it had a error in 7th bit,it would have been,

10101000

D3 D2 D1 D0 P3 P2 P1 P0
1 1 1 1 1 1 1 1
1 1 1 0 0 1 1 0
1 1 0 1 0 0 1 0
1 0 1 1 0 0 0 1

here we can see P1 value changes